OpenStudy (anonymous):

Simplify: (r-4) over (r^2-r-6) divided by (4-r) over (r+2) I know the answer is -1/(r-3), but I can't figure out how to get it there.

OpenStudy (callisto):

Factor (r^2-r-6) first.. (r^2-r-6) = r^2 -3r + 2r - 6 = r ( r-3) + 2 (r-3) = ...?

OpenStudy (anonymous):

When I factored (r^2-r-6) I got (r+2)(r-3)

OpenStudy (anonymous):

so then, do I multiply both sides of the equation by (r+2)(r-3)over 1?

OpenStudy (callisto):

(r-4) over (r^2-r-6) divided by (4-r) over (r+2) \[\frac{(r-4)}{(r^2-r-6)} \div \frac{(4-r)}{(r+2)}\]\[=\frac{(r-4)}{(r+2)(r-3)} \div \frac{(4-r)}{(r+2)}\]\[=\frac{(r-4)}{(r+2)(r-3)} \times \frac{(r+2)}{(4-r)}\]\[=\frac{(r-4)}{(r+2)(r-3)} \times \frac{(r+2)}{-(r-4)}\]

OpenStudy (callisto):

It's not an equation...

OpenStudy (anonymous):

sry, wrong terminology. not equation, problem

OpenStudy (callisto):

No worries... Do you get it?

OpenStudy (anonymous):

so then at the last step, do you just multiply straight across? then re-factor out?

OpenStudy (callisto):

Cancel the common factors.

OpenStudy (anonymous):

or wait, the (r+2)'s and the (r-4)'s cancel out.

OpenStudy (anonymous):

thank you so much. you were very helpful

OpenStudy (callisto):

Welcome :)